# High-energy physics image classification: A survey of Jet applications

### v1

@@ Abstract
In recent times, the fields of high-energy physics (HEP) experimentation and phenomenological studies have seen the integration of machine learning (ML) and its specialized branch, deep learning (DL). This survey offers a comprehensive assessment of these applications within the realm of various DL approaches. The initial segment of the paper introduces the fundamentals encompassing diverse particle physics types and establishes criteria for evaluating particle physics in tandem with learning models. Following this, a comprehensive taxonomy is presented for representing HEP images, encompassing accessible datasets, intricate details of preprocessing techniques, and methods of feature extraction and selection. Subsequently, the focus shifts to an exploration of available artificial intelligence (AI) models tailored to HEP images, along with a concentrated examination of HEP image classification pertaining to Jet particles. Within this review, a profound investigation is undertaken into distinct ML and DL proposed state-of-the art (SOTA) techniques, underscoring their implications for HEP inquiries. The discussion delves into specific applications in substantial detail, including Jet tagging, Jet tracking, particle classification, and more. The survey culminates with an analysis concerning the present status of HEP grounded in DL methodologies, encompassing inherent challenges and prospective avenues for future research endeavors.
@@ 1 Introduction
HEP, also known as particle physics, is a fascinating and complex branch of science that delves into the fundamental constituents of the universe and the forces that govern their interactions. This field explores the smallest building blocks of matter and the nature of the physical laws that govern their behavior at incredibly high energies. At its core, HEP aims to understand the most basic elements of matter and the fundamental forces that shape the universe. It seeks to answer questions about the origins of the universe, the nature of dark matter and dark energy, the unification of fundamental forces, and the properties of particles that make up matter and energy. Researchers in this field utilize sophisticated tools and massive particle accelerators, such as the large hadron collider (LHC), to probe matter at energy scales that are otherwise inaccessible. These colossal machines, which play a central role in HEP, accelerate subatomic particles to nearly the speed of light before colliding them, generating energy densities similar to those that existed just moments after the Big Bang. The collisions produced in these accelerators allow scientists to observe and study exotic particles that are often short-lived and can provide insights into the underlying structure of the universe.
The standard model (SM) of particle physics, which is the current theoretical framework in the field, describes the known elementary particles and their interactions. However, many mysteries remain unsolved, including the nature of dark matter, the absence of certain symmetries in the universe, and the lack of a unifying theory that reconciles gravity with the other fundamental forces. HEP has practical applications in technology and medicine, techniques and technologies developed in this field have led to advancements in medical imaging, radiation therapy, and materials science.
@@
The data acquisition system of LHC stores the data on tape using grid computing, it can be disseminated for offline analysis aimed at extracting information concerning particle trajectories formed within the detectors. These trajectories contain concealed details about numerous particle characteristics. Computer vision techniques become relevant and play a crucial role during the analysis of offline data. Specifically, in the realm of HEP data analysis, ML algorithms have found success, leading to significant enhancements in event classification performance when contrasted with traditional methods rooted in expert understanding. Techniques like boosted decision trees (BDT), shallow neural networks, and similar approaches have been employed in HEP data analysis. More recently, deep neural network (DNN) or DL have gained widespread adoption due to their applicability to intricate data structures such as images, videos, natural language, or sensor data. There are ongoing investigations into applying DNNs for analyzing granular details like particle positions and momentum as they traverse the detector. This has shown increased effectiveness in selecting signal events compared to ML algorithms employing conventional feature variables rooted in physics knowledge [1].
@@ 1.1 Motivation
In HEP, a track typically refers to the trajectory or path followed by a charged particle as it moves through a particle detector. HEP experiments often involve the collision of high-energy particles, such as those produced in particle accelerators like LHC. When these particles collide, they produce various other particles as a result of the interaction. These newly created particles then pass through a series of detectors designed to measure their properties. Each charged particle leaves behind a trace or track as it interacts with the detector’s various components, such as tracking chambers or silicon detectors. These tracks provide information about the particle’s momentum, charge, and the path it took through the detector. Analyzing these tracks is crucial for understanding the physics of the collisions and for identifying the types of particles produced.
The reconstruction of particle tracks involves sophisticated algorithms and software that piece together the recorded data from various detector components to reconstruct the paths of the particles accurately. These reconstructed tracks are essential for a wide range of analyses in HEP, including the discovery of new particles, the measurement of particle properties, and the investigation of fundamental forces and interactions in the universe.
However, in HEP experiments, there are always chances for high background signals or events that are not of primary interest to the experiment but can interfere with the detection and measurement of the particles or phenomena under investigation. The background sources could be the electronic components in the detector systems, when high-energy particles pass through matter, they can generate secondary particles through interactions, and others.
In the shade of the aforementioned phenomena and challenges, treating tracks / Jets in HEP as image-like data for processing and analysis is a useful approach, especially when dealing with the output from particle detectors. Hence, ML and DL play vital roles in HEP experiments. They serve the following purposes: i) Identifying and classifying particles by analyzing their tracks and energy deposits in detectors, thereby enhancing precision and identification speed. ii) Assisting in the accurate reconstruction of particle tracks from detector data, particularly in complex environments with numerous particles and interactions. iii) Enabling efficient data analysis, aiding researchers in sifting through extensive datasets to pinpoint rare or noteworthy events or particles. iv) Detecting anomalies or unexpected patterns in data, which could potentially signify the existence of new particles or physics beyond the SM, among other applications. These contributions underscore the significance of ML and DL in advancing HEP research.
@@ 1.2 Related work
In recent years, there has been a surge in reviews addressing various aspects of HEP [2, 3, 4, 5, 6]. Review [2] delved into the realm of supervised deep learning applied to high-energy phenomenology, discussing specific use cases such as employing machine learning to explore new physics parameter spaces and utilizing graph neural networks for particle production and energy measurements at the LHC. Meanwhile, paper [3] provided an overview of the initial forays into quantum machine learning in the context of HEP and offered insights into potential future applications. In [4], an array of novel tools relevant to HEP were introduced, complete with assessments of their performance, though there was limited discussion about future prospects. Lastly, the reviews [5, 6] comprehensively examined both theoretical and experimental aspects of Jets such as triggering, data acquisition systems, propagation, interactions, and related phenomena in HEP.
Table 1 provided assesses how the proposed review aligns with previous research in the field of HEP. Based on the assessment, it appears that our proposed review aims to comprehensively cover a wide range of topics in the field of ML and DL-based HEP, including Jet preliminaries, taxonomy of HEP, available Jet datasets, Jet image preprocessing, quantum machine learning, DL models for Jet images, classification techniques, Jet image DL applications, and research gaps/future directions. This suggests that the proposed review aims to provide a comprehensive overview of the current state of research in HEP and potential avenues for future work.
@@
Reference
Paper type
Publication year
Jet Preliminaries
Taxonomy of HEP
Available Jet datasets
Jet images pre-proce.
Quantium ML for
Jet images
DL models for Jet
images
Jet image DL classif.
techniques
Jet image DL apps
Research gaps and
future direction
[2]
Mini review
2019
[3]
Review
2021
[4]
Review
2021
[5]
Review
2022
[7]
Review
2022
[6]
Review
2023
Our
Review
2024
@@ 1.3 Contribution and survey structure
The objective of this survey is to provide a robust foundation for both HEP researchers aiming to grasp the principles of deep learning and its applications within the HEP domain, and for computer science researchers familiar with AI seeking insights into the fundamental features and prerequisites essential for constructing a robust AI model tailored specifically for HEP, employing Jet images. To achieve this goal, our contribution is encapsulated in the following key points:
@@
  • Different state-of-the-art methods are clustered based on the AI techniques employed and comprehensively reviewed accordingly. Additionally, the exploration of AI-based applications in HEP Jet images is thoroughly detailed.
@@
Figure 1: Mind map of the proposed review.
@@ 2.1 Types of particles
W and Z bosons: are important closely related particles described by the SM of particle physics. They and plays a significant role in the weak nuclear force, which is responsible for certain types of particle interactions and radioactive decay. The existence and properties of the $Z$ boson, along with the $W$ bosons, provided strong support for the electroweak theory and the SM as a whole. However, as with the $W$ boson, the Standard Model has limitations and does not explain all aspects of particle physics, such as gravity, dark matter, and the hierarchy of particle masses. Here are some key points about the W and Z bosons:
@@
The Higgs boson: is crucial to our understanding of how other particles acquire mass and, by extension, how the universe’s structure and behavior arise. The key points about the Higgs boson are:
@@
The top quark: is one of the fundamental particles described by the Standard Model of particle physics. It holds a special place in particle physics due to its extremely large mass and its role in various processes involving high-energy collisions. Here are some key points about the top quark:
@@ 2.2 Performance measures
In the realm of HEP, performance assessment is divided into two main categories. The first encompasses classical metrics like energy loss, path length, and axis distance. The second involves metrics utilized are related to DL-based HEP techniques, such as accuracy, true positive rate (TPR), false-positive rate (FPR), receiver operating characteristic (ROC), area under curve (AUC), mean squared error (MSE), Fubini-study tensor (FST), among others. Table LABEL:tab:metrics outlines these metrics, including mathematical formulations and descriptions.
Metric
Formula
Description
Classic metrics
Energy loss ratio
$\chi_{jh}\equiv\frac{E_{f}^{h}}{E_{i}^{h}}$
determine, on a Jet-by-Jet basis, the amount of energy loss.The notation "jh" denotes the energy of the Jet at the hadronic scale. Here, $E_{f}^{h}$ represents the $p_{T}$ of a specific Jet when interacting with a medium, whereas $E_{i}^{h}$ indicates the $p_{T}$ of the identical Jet in the absence of any medium.
Traversed path-length
$L=\frac{\sum_{i\in Jet}p_{T,i}L_{i}}{\sum_{i\in Jet}p_{T,i}}$
Offers significant insights closely related to the alterations and energy dissipation encountered by a Jet. This involves calculating the path length, $L$, traversed by a parton Jet, which is derived from the sum of the lengths of the Jet’s constituents at the partonic level $L_{i}$, weighted by their $p_{T}$.
Jet axis distance
$\Delta R=\sqrt{(\Delta\eta)^{2}+(\Delta\phi)^{2}}$
The constituents of a Jet are predominantly confined within a conical shape, characterized by a separation from the Jet axis of $\Delta R=0.5$. This separation $\Delta R$ encompasses both $\Delta\eta$, the variation in pseudo-rapidity, and $\Delta\phi$, the deviation in azimuthal angle relative to the Jet axis.
@@
Eff = $\frac{N_{top}^{top}}{N_{top}},\quad\text{Mistag}=\frac{N_{QCD}^{top}}{N_{QCD}}$
Defining efficiency and mis-tag rates is crucial for assessing the performance of the Artificial Neural Network (ANN) tagger. Where $N_{\text{top}}$ and $N_{\text{QCD}}$ represent the total jet count in the top and QCD jet samples, respectively, and $N_{top}^{QCD}$ signifies the number of jets in sample $top$ tagged as type $QCD$.
FPR and TPR
$\mathrm{\frac{FP}{FP+TN}},\mathrm{\frac{TP}{TP+FN}}$
The FPR, is the ratio (or percentage) of the background signal that are incorrectly identified as containing Jet. The TPR. is the ratio (or percentage) of the Jet signal that is correctly identified as Jet (particle).
AUC
$\int_{0}^{1}\mathrm{TPR}\>d(\mathrm{FPR})$
The area beneath the ROC curve is represented. It delivers a singular numeric score reflecting the cumulative effectiveness of the classification technique. An elevated AUC score signifies superior performance, with the ideal score being 1.
Accuracy
$\mathrm{\frac{TP+TN}{TN+FN+TP+FP}}$
The accuracy is the ratio (or percentage) of correctly detected instances of Jet in the signal. A high accuracy indicates that the classification algorithm is more effective in detecting Jet than background.
Deep learning metrics
MSE
$E(\theta)=\frac{1}{N}\sum_{i=1}^{N}(P_{b}^{i}(\theta)-T^{i})^{2}$
The training procedure seeks to discover the model parameter values denoted as $\theta$, which minimize the loss function known as MSE. Where $N$ is the number of training Jets, $P_{b}^{i}$ and $T^{i}$ is the predicted and target probabilities, respectively, for the $i$-th Jet [8].
F-measure
$2.\mathrm{\frac{Precision\times Recall}{Precision+Recall}}$
Represents the harmonic mean between precision and recall metrics. This measure is applied to assess the comprehensive efficacy of the classification algorithm in identifying or tagging Jets.
Joint and conditional distributions
$P(v\backslash O)\neq P(v)\hskip 8.5359pt\forall O$
If the probability distribution of the physics variable $v$, given the neural network output $O$, is influenced by the neural network output itself, then we interpret this as an indication that the network has acquired knowledge about this physics characteristic.
FST
$\frac{1}{{(1+\sum_{0}^{n}|z_{n}|^{2})^{2}}}\sum_{i=0}^{n}|dz_{i}|^{2}$
The FST is an invariant metric tensor that can be used to describe distances between quantum states [10, 11]. $z_{i}$ are complex coordinates, and $|dz_{i}|$ represents the absolute value of the differential of $z_{i}$.
AMS
$\sqrt{2\left((S+B+B_{r})\ln\left(1+\frac{S}{B+B_{r}}\right)-S\right)}$
The effectiveness of a method in categorizing test data into signal or background groups is assessed by employing approximate median significance (AMS) metric [4]. $S$ is the sum of weights of true positives; $B$ is the sum of weights of false positives; $B_{r}$ is the regularization constant.

### v2

# Image Classification in High-Energy Physics: A Comprehensive Survey of Applications to Jet Analysis

@@ Abstract
Nowadays, there has been a growing trend in the fields of high-energy physics (HEP) in its both parts experimental and phenomenological studies, to incorporate machine learning (ML) and its specialized branch, deep learning (DL). This review paper provides a thorough illustration of these applications using different DL approaches. The first part of the paper examines the basics of various particle physics types and sets up guidelines for assessing particle physics alongside the available learning models. Next, a detailed classification is provided for representing the jet images that are reconstructed in high energy collisions mainly with proton-proton collisions at well defined beam energies, covering various datasets, preprocessing techniques, and feature extraction and selection methods. The presented techniques can be applied to future hadron-hadron colliders (HLC) such as high luminosity LHC (HL-HLC) and future circular collider - hadron-hadron (FCC-hh). Next, the authors explore a number of AI models analysis designed specifically for images in HEP. We additionally undertake a closer look at the classification associated with images in hadron collisions, with an emphasis on Jets. In this review, we look into various state-of-the-art (SOTA) techniques in ML and DL, examining their implications for HEP demands. More precisely, this discussion tackles various applications in extensive detail, such as Jet tagging, Jet tracking, particle classification, and more. The review concludes with an analysis of the current state of HEP, using DL methodologies. It covers the challenges and potential areas for future research that will be illustrated for each application.
@@ 1 Introduction
High-energy physics (HEP) is an attracting and delicate branch of physics that manifests at the microscopic scale and which explores the fundamental building blocks of the universe and forces that govern their interactions at incredibly high energies under extremely intense conditions [1, 2]. In this field many sophisticated instruments and tools with large particle accelerators, like the current CERN-LHC (located near the French and Swiss border), to study matter at energy levels that are otherwise unattainable to reach with conventional methods. These gigantic machines accelerate subatomic particles at nearly the speed of light and then smash them together, creating energy densities analogous of the early moments after the Big Bang [3, 4]. By studying the collisions generated in these accelerators setups, it could be possible to track and evaluate rare particles that have a very short life time. This important study with the accumulated big data at higher collider luminosity values offers an improved understanding of the basic anatomy of different physics process and their topologies [5, 6].
The standard model (SM) is the present theoretical framework that describes the elementary particles and their interactions [7]. Despite its tremendous success in explaining many phenomena in nature, several mysteries remain unsolved, such as matter antimatter asymmetry, the nature of Dark Matter(DM), the neutrino mass and the hierarchy problem and many other open questions [8]. Furthermore, it is worth noting that besides its deep investigation about the Universe puzzles, HEP has demonstrated significant practical utility when used with advanced technologies [9]. As a matter of fact, the development of many techniques and technologies in this sector has driven notable progress in medical imaging [10, 11], radiation therapy [12, 13], and materials research [14, 15].
@@
The data acquisition system of large hadron collider (LHC) stores the data on tape using grid computing facilities, it can be disseminated for offline analysis aimed at extracting information concerning particle trajectories formed within the detectors. These trajectories contain concealed details about numerous particle characteristics. Computer vision techniques become relevant and play a crucial role during the analysis of offline data. Specifically, in the realm of HEP data analysis, machine learning (ML) algorithms have found success, leading to significant enhancements in event classification performance when contrasted with traditional methods rooted in expert understanding. Techniques like boosted decision trees (BDT), shallow neural networks, and similar approaches have been employed in HEP data analysis. More recently, deep neural network (DNN) or deep learning (DL) have gained widespread adoption due to their applicability to intricate data structures such as images, videos, natural language, or sensor data. There are ongoing investigations into applying DNNs for analyzing granular details like particle positions and momentum as they traverse the detector. This has shown increased effectiveness in selecting signal events compared to ML algorithms employing conventional feature variables rooted in physics knowledge [16].
@@ 1.1 Motivation
In HEP, a track typically refers to the trajectory or path followed by a charged particle as it moves through a particle detector. HEP experiments often involve the collision of high-energy particles, such as those produced in particle accelerators like LHC. When these particles collide, they produce various other particles as a result of the collisions. These newly created particles then pass through several sub-detectors where each designed to measure their corresponding properties. Each charged particle leaves behind a trace or track as it interacts with the detector’s various components, such as tracking chambers or silicon detectors. These tracks provide information about the particle’s momentum, charge, and the path it took through the detector. Analyzing these tracks is very crucial for understanding the physics of the collisions and for identifying the types of particles produced.
The reconstruction of particle tracks involves sophisticated algorithms and software that piece together the recorded data from various detector components to reconstruct the paths of the particles accurately. Then, the reconstructed tracks are essential for a wide range of analyses in HEP, including the discovery of new particles, the measurement of particle properties, and the investigation of fundamental forces and interactions in the universe.
However, in HEP experiments, there are always chances for high background contributions or events that are not of primary interest and can eventually mimic the physics signal and moreover can interfere along the physics collision. The background sources could be the electronic components in the different detector systems, when high-energetic particles pass through the material budget of the detector, they can also generate secondary tracks through different interactions, and possible decay modes.
In the light of the aforementioned phenomena and challenges, treating tracks / Jets in HEP as image-like data for processing and analysis is a useful approach, especially when dealing with the output from particle detectors. Hence, ML and DL play vital roles in HEP experiments. They serve the following purposes: i) Identifying and classifying particles by analyzing their tracks and energy deposits in detectors, thereby enhancing precision and identification speed, ii) assisting in the accurate reconstruction of particle tracks from detector data, particularly in complex environments with numerous particles and interactions, iii) enabling efficient data analysis schemes, one can sift through extensive datasets to pinpoint rare or noteworthy events or particles, iv) detecting anomalies or unexpected patterns in the recorded data, which could potentially signify the existence of new particles and physics beyond the SM, among other applications. These contributions underscore the significance of ML and DL in advancing HEP research topics.
@@ 1.2 Related work
In recent years, there has been a surge in reviews addressing various aspects of HEP [17, 18, 19, 20, 21]. The review presented in [17] delved into the realm of supervised deep learning applied to high-energy phenomenology, discussing specific use cases such as employing ML to explore new physics parameter spaces and utilizing graph neural networks for particle production and energy measurements at the LHC. Meanwhile, paper [18] provided an overview of the initial forays into quantum ML in the context of HEP and offered insights into potential future applications. In [19], an array of novel tools relevant to HEP were introduced, complete with assessments of their performance, though there was limited discussion about future prospects. Lastly, the reviews [20, 21] comprehensively examined both theoretical and experimental aspects of Jets such as triggering, data acquisition systems, propagation, interactions, and related phenomena in HEP.
Table 1 assesses how the proposed review aligns with previous research in the field of HEP. Based on the assessment, it appears that our proposed review aims to comprehensively cover a wide range of topics related to the ML and DL-based in HEP, including Jet preliminaries, taxonomy of HEP, available Jet datasets, Jet image preprocessing, quantum ML, DL models for Jet images, classification techniques, Jet image DL applications, and research gaps/future directions. This suggests that the proposed review aims to provide a comprehensive overview of the current state of research in HEP and potential avenues for future work.
@@
Reference Paper type Publication year Jet prelimanaries Taxonomy of HEP Available Jet datasets Jet images pre-process Quantium ML for Jet images DL models for Jet images Jet image DL classif. techniques Jet image DL apps Research gaps and future direction
[17] Mini review 2019
[18] Review 2021
[19] Review 2021
[20] Review 2022
[22] Review 2022
[21] Review 2023
Our Review 2024
@@ 1.3 Contribution and survey structure
The objective of this survey is to provide a robust foundation for both HEP researchers aiming to grasp the principles of deep learning and its applications within the HEP domain, and for computer science researchers familiar with artificial intelligence (AI) seeking insights into the fundamental features and prerequisites essential for constructing a robust AI model tailored specifically for HEP, employing Jet images. To achieve this goal, our contribution is encapsulated in the following key points:
@@
  • Different state-of-the art (SOTA) methods are clustered based on the AI techniques employed and comprehensively reviewed accordingly. Additionally, the exploration of AI-based applications in HEP Jet images is thoroughly detailed.
@@
Figure 1: Mind-map of the proposed review.
@@ 2.1 Types of particles
W and Z bosons: are important closely related particles described by the SM of particle physics. They are together known as the weak bosons or more generally as the intermediate vector bosons and plays a significant role in the weak nuclear force, which is responsible for certain types of specific interactions and radioactive decay. The existence and properties of the $Z$ boson, along with the $W$ bosons, provided strong support for the electroweak theory and the SM as a whole. However, as with the $W$ boson, the Standard Model has limitations and does not explain all aspects of particle physics, such as gravity, dark matter, and the hierarchy of particle masses. Here are some key points about the $W$ and $Z$ bosons:
@@
The Higgs boson: is crucial to our understanding of how other particles acquire mass and, by extension, how the universe’s structure and behavior arise. The key points about the Higgs boson are [23]:
@@
The top quark: is one of the heavy fundamental particles described by the Standard Model. It holds a special place in particle physics due to its extremely large mass and its role in various processes involving high-energy collisions. Here are some key points about the top quark [24]:
@@ 2.2 Performance measures
In the realm of HEP, performance assessment is divided into two main categories. The first encompasses classical metrics like energy loss, path length, and axis distance. The second involves metrics related to DL-based HEP techniques, such as accuracy, true positive rate (TPR), false-positive rate (FPR), receiver operating characteristic (ROC), area under curve (AUC), mean squared error (MSE), Fubini-study tensor (FST), among others. Table 2 outlines these metrics, including mathematical formulations and descriptions.
Table 2: An overview of the metrics employed to evaluate performance in particle physics.
Metric Formula Description
Classic metrics Energy loss ratio $\chi_{jh}\equiv\frac{E_{f}^{h}}{E_{i}^{h}}$ Determine, on a Jet-by-Jet basis, the amount of energy loss.The notation "jh" denotes the energy of the Jet at the hadronic scale. Here, $E_{f}^{h}$ represents the $p_{T}$ of a specific Jet when interacting with a medium, whereas $E_{i}^{h}$ indicates the $p_{T}$ of the identical Jet in the absence of any medium.
Traversed path-length $L=\frac{\sum_{i\in Jet}p_{T,i}L_{i}}{\sum_{i\in Jet}p_{T,i}}$ Offers significant insights closely related to the alterations and energy dissipation encountered by a Jet. This involves calculating the path length, $L$, traversed by a parton Jet, which is derived from the sum of the lengths of the Jet’s constituents at the partonic level $L_{i}$, weighted by their $p_{T}$.
Jet axis distance $\Delta R=\sqrt{(\Delta\eta)^{2}+(\Delta\phi)^{2}}$ The constituents of a Jet are predominantly confined within a conical shape, characterized by a separation from the Jet axis of $\Delta R=0.5$. This separation $\Delta R$ encompasses both $\Delta\eta$, the variation in pseudo-rapidity, and $\Delta\phi$, the deviation in azimuthal angle relative to the Jet axis.
@@
Eff = $\frac{N_{top}^{top}}{N_{top}},\quad\text{Mistag}=\frac{N_{QCD}^{top}}{N_{QCD}}$ Defining efficiency and mis-tag rates is crucial for assessing the performance of the artificial neural network (ANN) tagger. Where $N_{\text{top}}$ and $N_{\text{QCD}}$ represent the total jet count in the top and QCD jet samples, respectively, and $N_{top}^{QCD}$ signifies the number of jets in sample $top$ tagged as type $QCD$.
FPR and TPR $\mathrm{\frac{FP}{FP+TN}},\mathrm{\frac{TP}{TP+FN}}$ The FPR, is the ratio (or percentage) of the background signal that are incorrectly identified as containing Jet. The TPR. is the ratio (or percentage) of the Jet signal that is correctly identified as Jet (particle).
AUC $\int_{0}^{1}\mathrm{TPR}\>d(\mathrm{FPR})$ The area beneath the ROC curve is represented. It delivers a singular numeric score reflecting the cumulative effectiveness of the classification technique. An elevated AUC score signifies superior performance, with the ideal score being 1.
Deep learning metrics Accuracy $\mathrm{\frac{TP+TN}{TN+FN+TP+FP}}$ The accuracy is the ratio (or percentage) of correctly detected instances of Jet in the signal. A high accuracy indicates that the classification algorithm is more effective in detecting Jet than background.
MSE $E(\theta)=\frac{1}{N}\sum_{i=1}^{N}(P_{b}^{i}(\theta)-T^{i})^{2}$ The training procedure seeks to discover the model parameter values denoted as $\theta$, which minimize the loss function known as MSE. Where $N$ is the number of training Jets, $P_{b}^{i}$ and $T^{i}$ is the predicted and target probabilities, respectively, for the $i$-th Jet.
F-measure $2.\mathrm{\frac{Precision\times Recall}{Precision+Recall}}$ Represents the harmonic mean between precision and recall metrics. This measure is applied to assess the comprehensive efficacy of the classification algorithm in identifying or tagging Jets.
Joint and conditional distributions $P(v\backslash O)\neq P(v)\hskip 8.5359pt\forall O$ If the probability distribution of the physics variable $v$, given the neural network output $O$, is influenced by the neural network output itself, then we interpret this as an indication that the network has acquired knowledge about this physics characteristic.
FST $\frac{1}{{(1+\sum_{0}^{n}|z_{n}|^{2})^{2}}}\sum_{i=0}^{n}|dz_{i}|^{2}$ The FST is an invariant metric tensor that can be used to describe distances between quantum states [27, 28]. $z_{i}$ are complex coordinates, and $|dz_{i}|$ represents the absolute value of the differential of $z_{i}$.
AMS $\sqrt{2\left((S+B+B_{r})\ln\left(1+\frac{S}{B+B_{r}}\right)-S\right)}$ The effectiveness of a method in categorizing test data into signal or background groups is assessed by employing approximate median significance (AMS) metric [19]. $S$ is the sum of weights of true positives; $B$ is the sum of weights of false positives; $B_{r}$ is the regularization constant.